# Do we really have to filter out random noise in pre-training data for language models?

### v1

@@ Abstract
Web-scale pre-training datasets are the cornerstone of LLMs’ success. However, text data curated from the internet inevitably contains random noise caused by decoding errors or unregulated web content. In contrast to previous works that focus on low quality or synthetic data, our study provides the first systematic investigation into such random noise through a cohesive “What-Why-How” framework. Surprisingly, we observed that the resulting increase in next-token prediction (NTP) loss was significantly lower than the proportion of random noise. We provide a theoretical justification for this phenomenon, which also elucidates the success of multilingual models. On the other hand, experiments show that the model’s performance in downstream tasks is not based solely on the NTP loss, which means that random noise may result in degraded downstream performance. To address the potential adverse effects, we introduce a novel plug-and-play Local Gradient Matching loss, which explicitly enhances the denoising capability of the downstream task head by aligning the gradient of normal and perturbed features without requiring knowledge of the model’s parameters. Additional experiments on 8 language and 14 vision benchmarks further validate its effectiveness.
Do we really have to filter out random noise in pre-training data for language models?
Jinghan Ru, Yuxin Xie, Xianwei Zhuang, Yuguo Yin, Yuexian Zouthanks: Corresponding author. School of Electronic and Computer Engineering, Peking University {jinghanru, yuxinxie, xwzhuang, ygyin}@stu.pku.edu.cn zouyx@pku.edu.cn
@@ 1 Introduction
Figure 1: Overview of the study and methodology. (a) The common scenario in which a GPT model, pre-trained on filtered data $P^{c}$, demonstrates robust performance. (b) When the pre-training dataset is contaminated with random noise $P^{n}$, the resultant language model may exhibit unpredictable behavior. (c) Our approach focuses on the effective fine-tuning of black-box noisy models for downstream tasks $P^{d}$.
@@
Large language models (LLMs), particularly the GPT series (Radford et al., 2019; Brown, 2020; OpenAI, 2023), have fundamentally transformed the research landscape in natural language processing. The remarkable performance of these autoregressive models is largely attributed to pre-training on extensive datasets, which are gathered by crawling text from the whole internet. Given the sheer volume of these datasets, they inevitably encompass a wide variety of noise (Longpre et al., 2024; Elazar et al., 2024). Consequently, it is imperative to understand its impact, as the quality of pre-training data plays a decisive role in the effectiveness of LLMs (Touvron et al., 2023). Allen-Zhu and Li (2024a); Xie et al. (2023b) highlight that low-quality data can significantly diminish a model’s knowledge capacity and performance. Shumailov et al. (2024); Seddik et al. (2024) demonstrate that recursively training LLMs with synthetic data can lead to model collapse.
@@
However, little attention has been paid to the impact of random noise within datasets. Due to anti-crawling mechanisms (Gao et al., 2023), decoding errors [^1] https://stackoverflow.com/questions/62499600/gibberish-text-output-because-of-encoding-in-web-scraping, and tremendous amounts of unmaintained websites [^2] https://community.cloudflare.com/t/website-showing-garbage-text, the raw data obtained through web crawling inevitably contains a substantial amount of random noise (Zhou et al., 2024; Chen et al., 2022; Kang et al., 2023). Although theoretically it may not be challenging to remove such noise, practical limitations in computational resources often result in incomplete data cleaning(Albalak et al., 2024; Soldaini et al., 2024). For example, it is observed that the Chinese corpus used to train the GPT-4o tokenizer contains a considerable amount of nonsensical data.  [^3] https://github.com/jiangyy/gpt-tokens Therefore, it is of great importance to gain a thorough understanding of the effects of such random noise.
@@
We conduct extensive experiments based on the OpenWebText dataset (Gokaslan et al., 2019) used to pre-train language models with the same architecture and parameter size as GPT-2. Specifically, to simulate random noise shown in Figure 1, we randomly generate, with proportions of 1%, 5% and 20%, a sequence of integers within the range of 0 to 50256, according to the vocabulary size of GPT-2’s tokenizer, to simulate the tokenization outcome of nonsensical text found on the internet. Interestingly, we observe that the presence of random noise does not lead to a catastrophic failure in model training; instead, its effect on autoregressive loss is disproportionately small, e.g., the increase in loss is only about 1% even with 20% of the dataset being noisy. We provide a theoretical analysis to explain these phenomena, which also sheds light on the success of multilingual models (where one language may appear as “noise” to another) and large speech models (Chen et al., 2022), indicating the broader implications of studying the effects of random noise.
@@
On the other hand, further experiments reveal that a model that exhibits a lower NTP loss experiences a 1.5% decrease in accuracy on downstream tasks. This indicates that performance on downstream tasks is not solely rely upon the NTP loss. Given the common practice of fine-tuning a pre-trained foundation model rather than undergoing a full pre-training process from scratch, we opt to follow the work of Chen et al. (2024) by exploring how to efficiently fine-tune language models only using extracted features for downstream tasks when the pre-training data and model weights are not accessible, which reflects real-world application scenarios for LLMs. To mitigate the potential adverse effects of noise, we propose a novel plug-and-play Local Gradient Matching (LGM) loss. This method involves artificially adding noise to the output features and minimizing the gradient difference between the noisy and original features. We also provide a theoretical analysis of the LGM loss. Interestingly, when applying the LGM loss to fine-tune clean models such as Llama-3 or ViT-L on 8 language and 14 vision datasets, we observe an unexpected improvement in accuracy, which effectively demonstrates the versatility and broad applicability of the LGM loss beyond its original intent of addressing noise-related issues.
@@
  • What: What is the effect of random noise? In Section 3, we demonstrate through experiments how random noise impacts NTP loss.
@@
Section 4 further provides evidence of the novelty and effectiveness of the LGM loss by conducting extensive experiments on 22 downstream tasks.
In summary, our contributions are as follows: (1) We investigate the underexplored problem of random noise in pre-training datasets for language models. (2) We pre-train multiple GPT-2 models and the empirical results show that the influence of random noise on NTP loss is insignificant. Then we provide a theoretical analysis, extending our findings to other domains and thus highlighting the significance of this research direction. (3) We propose a novel blackbox fine-tuning LGM loss for downstream tasks, supported by comprehensive experimental and theoretical analysis that confirm its efficacy. Code, data, and model checkpoint weights are available at this repo.
@@ 2 Related Works
Figure 2: Next-token prediction loss on the clean OpenWebText validation set for GPT-2 models pre-trained on synthetic OpenWebText datasets with varying levels of random noise. (a) Trend of NTP loss as training proceeds. (b) Difference in NTP loss between the noisy and clean models after the same number of training iterations. (c) Difference in loss values after undergoing the same number of training iterations on clean OpenWebText data.
Pre-training Data Analysis for Language Model Training. Elazar et al. (2024) analyzed open-source datasets like The Pile and C4, uncovering significant amounts of low-quality content in these datasets. Allen-Zhu and Li (2024a); Seddik et al. (2024) highlighted the negative impact of such data on training. Despite these remarkable contributions, there remains a lack of understanding regarding the specific effects of random noise on language model performance. This paper aims to address this gap.
Noisy Model Learning. Our work draws significant inspiration from Noisy Model Learning (NML) proposed by Chen et al. (2024). In NML, the authors introduce noise into large datasets like ImageNet by randomly altering labels, then pre-train neural networks on these noisy datasets. The study reveals that moderate label noise enhances in-distribution (ID) sample classification, while out-of-distribution (OOD) performance deteriorates with increasing noise. This paper extends the concept of NML, presenting theoretical insights and methodologies that are applicable across multiple modalities and various problems.
Due to space limitations, the detailed related works are provided in Appendix C.
Figure 3: Validation experiments. (a) Loss trends on the random noise in the training set of the model trained on the dataset with 5% random noise. (b) Comparison of the loss between 5% random noise and Gaussian noise. (c) The loss difference on the clean OpenWebText validation set compared to the baseline for models trained on datasets with 5% random noise and 5% Gaussian noise, respectively.
@@ 3 Revealing the Effect of Random Noise in Language Model Pre-training
3.1 Experimental Design
Preliminary. Let $L$ denote the maximum context length of the language model and let $\mathcal{W}$ represent the model’s vocabulary with size $V=|\mathcal{W}|$. We define $\mathcal{X}$ as the set of all discrete sentences that the model can represent, where $\mathcal{X}=\cup_{i=1}^{L}\{0,1,\dots,V-1\}^{i}=\cup_{i=1}^{L}\mathcal{W}^{i}$ and $\{0,1,\dots,V-1\}^{i}$ represents prefixes of length $i$. For any discrete set $A$, let $\Delta_{A}$ denote the set of all probability distributions defined on $A$. Given that next-token prediction (NTP) is actually a classification task given the prefix, we define joint probability distributions $P^{c},P^{n},P^{m}\in\Delta_{\mathcal{X}\times\mathcal{W}}$ where $P^{c}$ represents the distribution of clean data, $P^{n}$ represents the distribution of noise data, and $P^{m}$ represents the distribution of the mixed pre-training dataset which contains both clean and noise data. Since the noisy dataset can be viewed as the concatenation of clean data and random noise, it can be formalized by the Huber contamination model (Fang et al., 2022) as follows:
$P^{m}=\alpha P^{n}+(1-\alpha)P^{c}\vspace{-1.5mm}$ (1)
@@ 3.1 Experimental Design
We use $\mathcal{H}$ to denote the hypothesis space (e.g., all possible parameter configurations given the transformer architecture ). Define $h:\mathcal{X}\rightarrow\mathbb{R}^{V}\in\mathcal{H}$ as the language model and $\boldsymbol{p}^{h}_{\cdot|x}(w)$ as the $w$-th component of the probability distribution induced by $h(x)$. The next-token prediction loss can be expressed as follows:
@@
Data setup. We utilize the OpenWebText dataset (Gokaslan et al., 2019) which comprises 8 billion tokens as an alternative to the original WebText dataset used to train GPT-2 124M models (Radford et al., 2019). Concretely, to mimic the unpredictable nature of garbage noise after it has been tokenized, we first generate a sequence of integers where each integer follows a uniform distribution over $[0,50256)$ (recall that 50256 is the vocabulary size of GPT-2’s tokenizer). Furthermore, we notice that beyond uniform distribution, certain tokens appeare with significantly higher frequency in real-world random noise. To reflect this phenomenon, approximately 100 tokens in $[0,50256)$ are randomly selected and their frequencies increase accordingly. The overall distribution of the data can be referred to in Figure 8. The generated noise is then added to the clean dataset such that $\alpha$ is 1%, 5%, and 20% respectively. Each synthetic noisy dataset is used to pre-train a GPT-2 model. We argue that given the vast diversity of websites on the internet, it is inevitable that datasets used for training large models will inadvertently include such garbled text as shown in Figure 1. This kind of noise is not just theoretical; it is a practical issue encountered during web crawling and data collection processes.
We set the context length $L$ to be 1024 and the batch size to be 640. All models are trained for 300,000 iterations. To evaluate the performance, the resulting model checkpoints are tested on the clean OpenWebText validation set, measuring the NTP loss for comparison. Further details regarding datasets and experimental parameters can be found in Appendix D.
3.2 Results
In Figure 2, we illustrate the evolution of the NTP loss throughout the training process. Although random noise has a negative effect on the model’s performance as expected, experimental results yield two intriguing insights:
(1) In contrast to the low-quality or synthetic data, the presence of random noise does not lead to training collapse, even when the noise level reaches 20%. While increasing training time on low quality or synthetic data typically degrades model performance (Allen-Zhu and Li, 2024a; Shumailov et al., 2024), extending the training duration continues to drive down the model’s loss in the case of random noise.
(2) The impact of random noise on the loss is disproportionately small. For instance, 5% of random noise only results in a 0.2% increase in the NTP loss. This discrepancy becomes even smaller if the noisy models are calibrated to match the number of training iterations with the baselines trained on clean datasets.

### v2

@@ Abstract
Web-scale pre-training datasets are the cornerstone of LLMs’ success. However, text data curated from the Internet inevitably contains random noise caused by decoding errors or unregulated web content. In contrast to previous works that focus on low quality or synthetic data, our study provides the first systematic investigation of such random noise through a cohesive “What-Why-How” framework. Surprisingly, we observed that the resulting increase in the loss of next-token prediction (NTP) was significantly lower than the proportion of random noise even when the model was scaled up to 2.7B. We provide a theoretical justification for this phenomenon, which also elucidates the success of multilingual models and can be applied to multimodal models. On the other hand, experiments show that the model’s performance in downstream tasks is not based solely on the NTP loss, which means that random noise may result in degraded downstream performance. To address the potential adverse effects, we introduce a novel plug-and-play Local Gradient Matching loss, which explicitly enhances the denoising capability of the downstream task head by aligning the gradient of normal and perturbed features without requiring knowledge of the model’s parameters. Additional experiments on 8 language and 14 vision benchmarks further validate its effectiveness.
@@ 1 Introduction
Large language models (LLMs) have fundamentally transformed the research landscape in natural language processing. The remarkable performance of these autoregressive models is largely attributed to pre-training on extensive datasets, which are gathered by crawling text from the whole Internet. Given the sheer volume of these datasets, they inevitably encompass a wide variety of noise [1, 2]. Consequently, it is imperative to understand its impact, as the quality of pre-training data plays a decisive role in the effectiveness of LLMs [3]. Previous studies [4, 5] highlight that low quality data can significantly decrease the knowledge capacity and performance of a model, and recursive training of LLMs with synthetic data can lead to model collapse [6, 7].
@@
However, little attention has been paid to the impact of random noise within datasets. Due to anti-crawling mechanisms, decoding errors, and huge amounts of unmaintained websites, the raw data obtained through web crawling inevitably contains a substantial amount of random noise [8, 9, 10]. Although theoretically it may not be challenging to remove such noise, practical limitations in computational resources often result in incomplete data cleaning [11, 12]. Therefore, it is of great importance to gain a thorough understanding of the effects of such random noise.
@@
We conduct extensive experiments based on the OpenWebText dataset [13] used to pre-train language models with the same architecture as GPT-2 and parameter size up to 2.7B. Specifically, to simulate random noise shown in Figure 1, we randomly generate a sequence of random integers within the range of 0 to 50256, the vocabulary size of GPT-2’s tokenizer, to simulate the tokenization outcome of non-sensical text found on the Internet. Interestingly, we observe that the presence of random noise does not lead to a catastrophic failure in model training; instead, its effect on autoregressive loss is disproportionately small, e.g., the increase in loss is only about 1% even with 20% of the dataset being noisy. We provide a theoretical analysis to explain these phenomena, which also sheds light on the success of multilingual models [14] and multimodal LLMs [15, 16], indicating the broader implications of studying the random noise.
Figure 1: Overview of the study and methodology. (a) The common scenario in which a GPT model, pre-trained on filtered data $P^{c}$, demonstrates robust performance. (b) When the pre-training dataset is contaminated with random noise $P^{n}$, the resultant language model may exhibit unpredictable behavior. (c) Our approach follows [17] and focuses on the effective fine-tuning of black-box noisy models for downstream tasks $P^{d}$.
@@
On the other hand, further experiments reveal that a noisy model that exhibits a lower NTP loss experiences a 1.5% decrease in accuracy on downstream tasks. This indicates that performance on downstream tasks is not solely rely upon the NTP loss. Given the common practice of fine-tuning a pre-trained foundation model rather than undergoing a full pre-training process from scratch, we opt to follow the work of [17] by exploring how to efficiently fine-tune language models only using extracted features for downstream tasks when the pre-training data and model weights are not accessible, which reflects real-world application scenarios for LLMs. To mitigate the potential adverse effects of noise, we propose a novel plug-and-play Local Gradient Matching (LGM) loss. This method involves artificially adding noise to the output features and minimizing the gradient difference between the noisy and original features. We also provide a theoretical analysis of the LGM loss. Interestingly, when applying the LGM loss to fine-tune clean models such as Llama-3 or ViT-L on 8 language and 14 vision datasets, we observe an unexpected improvement in accuracy, which effectively demonstrates the versatility and broad applicability of the LGM loss beyond its original intent of addressing noise-related issues.
@@
  • What: What is the effect of random noise? In Section 3, we demonstrate through experiments how random noise affects NTP loss.
@@
Section 4 further provides evidence of the novelty and effectiveness of LGM loss by conducting extensive experiments on 22 downstream tasks.
Our contributions are as follows: (1) We investigate the underexplored problem of random noise in pre-training datasets for language models. (2) We pre-train multiple GPT-2 models with parameter size up to 2.7B and the empirical results show that its influence on NTP loss is insignificant. Then we provide a theoretical analysis that can be extended to other domains. (3) We propose a novel blackbox fine-tuning LGM loss for downstream tasks, supported by comprehensive experimental and theoretical analysis. Code, data, and model checkpoint weights are available at this repo.
@@ 2 Related Works
Pre-training Data Analysis for Language Model Training. Elazar et.al [2] analyzed open-source datasets like The Pile and C4, uncovering significant amounts of low-quality content in these datasets. Exsiting works [4, 7] highlighted the negative impact of such data on training. Despite these remarkable contributions, there remains a lack of understanding regarding the specific effects of random noise on language model performance. This paper aims to address this gap.
Noisy Model Learning. Our work draws significant inspiration from Noisy Model Learning (NML) [17]. In NML, the authors introduce noise into large datasets like ImageNet by randomly altering labels, then pre-train neural networks on these noisy datasets. The study reveals that moderate label noise enhances in-distribution (ID) sample classification, while out-of-distribution (OOD) performance deteriorates with increasing noise. This paper extends the concept of NML, presenting new theoretical insights and methodologies.
Due to space limitations, detailed related works are provided in Appendix C.
@@ 3 Revealing the Effect of Random Noise in Language Model Pre-training
3.1 Preliminary
Let $L$ denote the maximum context length of the language model and let $\mathcal{W}$ represent the model’s vocabulary with size $V=|\mathcal{W}|$. We define $\mathcal{X}$ as the set of all discrete sentences that the model can represent, where $\mathcal{X}=\cup_{i=1}^{L}\{0,1,\dots,V-1\}^{i}=\cup_{i=1}^{L}\mathcal{W}^{i}$ and $\{0,1,\dots,V-1\}^{i}$ represents prefixes of length $i$. For any discrete set $A$, let $\Delta_{A}$ denote the set of all probability distributions defined on $A$. Given that next-token prediction (NTP) is actually a classification task given the prefix, we define joint probability distributions $P^{c},P^{n},P^{m}\in\Delta_{\mathcal{X}\times\mathcal{W}}$ where $P^{c}$ represents the distribution of clean data, $P^{n}$ represents the distribution of noise data, and $P^{m}$ represents the distribution of the mixed pre-training dataset. Since the noisy dataset can be viewed as the concatenation of clean data and random noise, it can be formalized by the Huber contamination model [18] as follows:
$P^{m}=\alpha P^{n}+(1-\alpha)P^{c}$ (1)
@@ 3.1 Preliminary
We use $\mathcal{H}$ to denote the hypothesis space (e.g., all possible parameters given the transformer architecture ). Define $h:\mathcal{X}\rightarrow\mathbb{R}^{V}\in\mathcal{H}$ as the language model and $\boldsymbol{p}^{h}_{\cdot|x}(w)$ as the $w$-th component of the probability distribution induced by $h(x)$. The NTP loss can be expressed as follows:
@@
3.2 Random Noise Data Generation
3.2.1 What is “Random Noise”?
Our study focuses on “random noise”, informally defined as any token sequence that appears nonsensical with respect to the distributions humans can read. In real-world text data crawled from the Internet, we encounter at least two types of random noise: (1) Pure random data with a uniform distribution: such noise may arise from hardware faults on backend servers, encrypted content, or website defense mechanisms. For example, Son et al. [19] observe that “several defense techniques inject random noise during website rendering to degrade the attack success rate.” (2) Decoding-error data with a distorted distribution: this noise can originate from decoding errors, the vast volume of unmaintained websites  [^1] https://community.cloudflare.com//t//website-showing-garbage-text, and even appears in Oracle’s official documentation  [^2] https://docs.oracle.com/cd//E19528-01//820-0890//WebSvr.html#wp35099. Although such data follows some distribution, it diverges drastically from clean data, and experiments show that a GPT-2 model trained on clean data exhibits significantly higher predictive entropy with these sequences.
Since neither of these noise types occurs in a normal dataset, in mathematical terms this implies that for any prefix $r$ sampled from $P^{n}_{X}$, the probability under the clean distribution $P^{c}_{X}(r)$ is zero. Thus, we make the following mild assumption as the formal definition of random noise:
Assumption 1.
$P^{c}$ and $P^{n}$ have disjoint support sets, i.e., $\mathrm{supp}(P^{c})\cap\mathrm{supp}(P^{n})=\emptyset$.
Previous studies have focused on synthetic or low-quality data, which is similar to clean data distributions but contain incorrect or substandard information. An imperfect analogy is to treat next-token prediction as an image classification task: low-quality data corresponds to label noise in image classification, whereas random noise represents a fundamental corruption of the image distribution itself. This distinction marks the clear novelty of our work relative to prior efforts.
3.2.2 Data Generation Process
Due to constrained computational resources, we are unable to collect a sufficiently large corpus of genuine random noise from live web crawls. Instead, we adopt a synthetic approach that faithfully emulates the two primary noise modalities mentioned above. We use OpenWebText [13] as the clean dataset $D_{c}$ comprising $|D_{c}|\approx 8$ billion tokens as an alternative to the original WebText dataset.
Uniform Random Noise. To mimic pure random corruption (e.g., hardware faults, encryption artifacts, or anti-crawling defenses [19]), we generate sequences of token IDs following uniform distribution as the random noise data $D_{n}$. Concretely, let $V=50256$ denote the tokenizer vocabulary size and $\alpha$ the desired noise proportion. We sample
$x_{i}\sim\mathrm{Uniform}\bigl{(}0,V-1\bigr{)}\,,\quad i=1,\dots,|D_{n}|,$ (3)
where $|D_{n}|$ is chosen such that $\alpha=\frac{|D_{n}|}{|D_{n}|+|D_{c}|}$. Our experiments are primarily based on $D_{n}$. In the following text, unless otherwise specified, the term "random noise" refers to this particular type.
Gaussian-Distributed Noise. To capture decoding errors and web-rendering artifacts that follow certain skewed distribution, we further sample Gaussian noise data $D_{g}$ from a truncated discrete Gaussian distribution to mimic a learnable noise. Let $\mu=(V-1)/2$ and $\sigma$ be a chosen standard deviation (e.g., $\sigma=500$). For $i=1,\dots,|D_{g}|$, we draw
$x_{i}\sim\mathrm{DiscreteGaussian}(\mu,\sigma)\quad\text{clipped to }[0,V-1].$ (4)
Input: Clean dataset $D_{C}$, noise ratio $\alpha$, maximum context length $L$, batch size $B$
 // Data generation process
 $D_{n}=\emptyset$
while $\frac{|D_{n}|}{|D_{n}|+|D_{c}|}\leq\alpha$ do
       sample $x_{i}$ according to Equation (3) // or (4)
        $D_{n}$ = concat$(D_{n},x_{i})$
end while
$D_{m}$=concat$(D_{c},D_{n})$
// Pretraining Stage
while training not converged do
       // sample a mini-batch
       ix = torch.randint($|D_{m}|$ $L$, ($B$))
       x = [$D_{m}$[i:i+$L$] for i in ix]
       y = [$D_{m}$[i+1:i+1+$L$] for i in ix]
       l = cross_entropy(x, y)
       backward_loss_and_update_optimizer()
end while
Algorithm 1 Synthetic Noise Generation
Upon acquiring the noisy data, we directly concatenate it to $D_{c}$. Since token sampling during training is completely random, the positional arrangement of noise data - whether pre-appended, appended, or interleaved with clean data - becomes statistically irrelevant in practice. The complete data generation and pre-training procedures are formally described in Algorithm 1. We set $\alpha$ to be 1%, 5%, and 20% for $D_{n}$, and 5% for $D_{g}$.
While our pioneering experiments employ a relatively simple simulation strategy, we argut that it nevertheless captures the essential characteristics of real-world random noise . By adopting Gaussian or random noise as an initial fully parameterizable testbed, we ensure analytical tractability and reproducibility, and it can readily be extended to more realistic data in future work.
3.3 Experiments
Figure 2: Next-token prediction loss on the clean OpenWebText validation set for GPT-2 124M models pre-trained on synthetic OpenWebText datasets with varying levels of random noise. (a) Trend of NTP loss as training proceeds. (b) Difference in NTP loss between the noisy and clean models after the same number of training iterations. (c) Difference in loss values after undergoing the same number of training iterations on clean OpenWebText data.
Figure 3: Validation experiments with the models trained with 5% random noise. (a) Loss trends on the OpenWebText validation set for GPT-2 2.7B model. (b) Comparison of the 124M training set loss between 5% random noise and Gaussian noise. (c) Loss difference on the validation set for 124M models trained on datasets with 5% random noise and 5% Gaussian noise, respectively.
Our work is based on nanoGPT. Specifically, we pretrain GPT-2 on clean or noisy dataset with two parameter configurations: The first adheres to the GPT-2 124M [20] architecture with 12 transformer layers, 12 attention heads, and 768-dimensional embeddings. The second follows the GPT-Neo 2.7B [21] specification comprising 32 layers, 20 attention heads, and 2560-dimensional embeddings. Constrained by computational resources, our primary experiments utilize the 124M variant, while a 2.7B model pre-trained on 5% random noise data is strategically deployed to validate the scalability of our findings. After pretraining, the resulting model checkpoints are tested on the clean OpenWebText validation set, measuring the NTP loss for comparison. Further details regarding datasets and experimental parameters can be found in Appendix D.
In Figure 2, we illustrate the evolution of the NTP loss throughout the training process. Although random noise has a negative effect on the model’s performance as expected, experimental results yield intriguing insights: the impact of random noise on the loss is disproportionately small. In fact, according to conventional statistical learning theory [22], the performance of machine learning models degrades as the distribution shift between training and test data increases. For example, Ben-David et al. [23] have shown that the loss on the test set can be upper bounded by the sum of training set loss and total variation distance between the training and test distributions, which reaches the maximum value when the two distributions have disjoint support set. In other words, previous theory predicts that random noise will lead to worst-case model performance. However, our experiments show the contradictory results. For instance, 5% of random noise only results in a 0.2% increase in the NTP loss. This discrepancy becomes even smaller if the noisy models are calibrated to match the number of training iterations with the baselines trained on clean datasets.